Convert each number to the specified base. Be sure to include the base when needed.
Express 2713(8) in base 10 ________

Answer by jim_thompson5910(35256)

2713 base 8 = 2*8^3 + 7*8^2 + 1*8^1 + 3*8^0 base 10

2713 base 8 = 2*512 + 7*64 + 1*8 + 3*1 base 10

2713 base 8 = 1024 + 448 + 8 + 3 base 10

2713 base 8 = 1483 base 10